Please help me solve this question:
"A father is four times as old as his son. In four years' time their combined age will be 48 years. How old are they now?"

Answer by stanbon(75887)

A father is four times as old as his son. In four years' time their combined age will be 48 years. How old are they now?"
----
Equations:
f = 4s
(f+4)+(s+4) = 48
-----------------------
Substitute for "f" and solve for "s":
(4s+4)+(s+4) = 48
5s+8 = 48
5s = 40
s = 8 yrs old (son's age now)
--------------
Solve for "f":
f = 4s
f = 4*8 = 32 yrs old (father's age now)
==============================================
Cheers,
Stan H.
=========